OpenStudy (anonymous):

solve elimination process x+9y=85 -x+9y=77

OpenStudy (anonymous):

gw2011 no one helped me on this one solve elimination process x+9y=85 -x+9y=77

OpenStudy (gw2011):

x+9y=85 -x+9y=77 By adding these two equations you get: 18y=162 y=9 Substitute this value into one of the equations and you get: x+9y=85 x+(9)(9)=85 x+81=85 x=4 Therefore, x=4 and y=9

OpenStudy (anonymous):

thanks you did it again you are too smart

OpenStudy (gw2011):

Thanks and you're welcome